# Probing Langmuir monolayer self-assembly in condensed and collapsed phases: grazing incidence X-ray diffraction and X-ray standing waves studies

### v1

@@ Abstract
Ce-induced effects on the self-assembly of arachidic acid Langmuir monolayers was studied in this work. The monolayers were formed on the liquid subphase in the presence of Ce(III) ions. A new type of structural configuration is found for such monolayers, in which the monolayer maintains its structural ordering despite being compressed beyond the collapse point. Instead of forming 3D aggregates as in the typical collapsed state, the monolayer appears to be corrugated. Grazing incidence X-ray diffraction and X-ray standing waves confirm these findings. The diffraction pattern for the monolayer in a new state is represented by the unclosed diffraction rings with maxima near the sample horizon. This diffraction pattern is quantitatively reproduced in the numerical simulations by assuming the corrugated monolayer. The details of the conditions under which these corrugated Langmuir monolayers were observed and the analysis of the diffraction data are described.
@@ 1 Introduction
Scientific interest in the self-assembly of amphiphilic trivalent lanthanide complexes is motivated by their photophysical properties, which can be exploited for molecular sensing [1]. The practical realization of such systems requires the assembly of the lanthanide complexes into nanostructures [2]. One of the possible ways to create functional nanomaterials with predefined molecular architecture is the Langmuir-Blodgett technique [3], which involves the self-assembly of amphiphiles at the air-liquid interface into a Langmuir monolayer and subsequent transfer to a solid substrate. Charged headgroups of amphiphilic molecules are known to interact with metal ions when present in the liquid subphase [4]. In this way, self-assembly defines the performance of a nanodevice. Self-assembly is driven not only by the sort of amphiphiles and the thermodynamic conditions under which the monolayer is formed [5], but can also be directed by interactions with ions in the subphase [6]. Therefore, further development of such functional devices requires extensive and detailed studies of the structural organization of amphiphilic trivalent lanthanide complexes.
@@
There are two known scenarios of how the interaction with metal ions affects the structure of the Langmuir monolayer, as reported in [7] using divalent metal ions as an example. The first, more common scenario is that the interaction between headgroups and metal ions leads to the formation of a disordered thin monolayer of metal ions underneath the organic monolayer, which affects the compressibility of the monolayer, resulting in structures similar to high-pressure solid-phase monolayers [7]. The second scenario is the formation of the ionic monolayer ordered into a superstructure commensurate with the organic monolayer, which was first reported in [8]. The study reports the formation of a Cd2+ superstructure underneath an arachidic acid monolayer. The structure has a cell commensurate with the organic monolayer structure. The transition between these two scenarios of interaction of the organic monolayer with metal ions may depend on the pH of the subphase and the ion concentration. In [9], the interaction of behenic acid with a variety of divalent ions was studied, and an ion concentration threshold was observed. Above this threshold, the ions formed a superlattice, and the ion layer was disordered below it. Later studies [10, 11] report on forming ordered monolayers of lanthanide ions. They tested a wide range of trivalent lanthanides. They found that the ordering depends not only on the molecules forming the Langmuir monolayer, but also on the atomic number of a lanthanide. In addition, the crystalline structure of the ionic monolayer can be commensurate or incommensurate with the lattice of the organic molecules.
@@
Another interesting aspect is the collapse of Langmuir monolayers in the presence of metal ions. The study in [12] reports the spontaneous collapse of fatty acid monolayers in the presence of Ca2+ ions. It has also been shown both experimentally [13] and theoretically [14] that the presence of Ca2+ ions changes the structure of the collapsed fatty acid monolayer. Even more interesting is that the collapsed structures show a high degree of order, namely the formation of inverted bi-layers. The formation of such inverted bi-layers is also reported for lanthanides [15]. Yet again, the structure of the collapsed layers is ion-specific.
@@
Thus, the interaction of fatty acids with lanthanide ions has shown complex and widely varying results. However, understanding the mechanisms of ion monolayer formation and its effect on collapse is relevant to the problems of functionalizing lanthanides into self-assembled nanostructures. This article focuses on the self-assembly of arachidic acid monolayers in the presence of Ce3+ ions and their structural reorganization while compressed beyond the collapse point. We have studied the structure of fatty acid monolayers interacting with metal ions at the air-liquid interface at different surface pressures and temperatures. As the primary tool for monitoring the monolayer structure, we used grazing incidence X-ray diffraction (GID) at the bending magnet synchrotron radiation source. This choice is due to the sensitivity of GID to both translational and orientational lateral ordering of the monolayers. In addition, simultaneously with GID experiments, we performed the X-ray standing wave (XSW) measurements to monitor the vertical localization of Ce3+ ions during the structural reorganization at the collapse.
@@ 2.1 Surface pressure-area isotherms
The effect of Ce3+ ions on the thermodynamic properties of the arachidic acid monolayer was investigated in our additional series of experiments using compression isotherm measurements. We have studied the behavior of the arachidic acid monolayer, formed on Ce3+ solution and on pure water as a control. All measurements were performed at a temperature of C. The typical - isotherms are shown in Fig. 1. Based on these experimental results, the key feature of the arachidic acid monolayer formed in the presence of Ce3+ is the absence of a liquid phase, as evidenced by a high monolayer compressibility – 2.2 m/mN.
Our observations are consistent with the results reported in [16], where detailed studies of the binding interaction of lanthanide ions (La3+ and Y3+) with arachidic acid monolayers at submicromolar salt concentrations were reported. Note that similar changes in the isotherms were observed in [17], where the effect of Cu2+ on the organization and morphology of the arachidic acid monolayer was studied.
Generally, metal binding with carboxylate acid groups is recognized to promote the formation of a rigid monolayer with higher crystallinity and reduced molecular area [18]. The interaction of divalent and trivalent metallic ions with carboxylic groups of fatty acids is strongly dependent on such parameters as solution pH, temperature, salt concentration. The type of metal-carboxylate complex (ionic, unidentate, chelating bridging bidentate, polymeric tridentate) is determined to a large extent by metallic ion specific nature, i.e. ion size, coordination structure, oxidation state in solution, etc. Large Ln3+ ions with high coordination numbers exhibit rather intricate binding behavior [16], resulting in a wide variety of the lanthanide-carboxylate binding interactions. According to the data reported in [16], at least three different coordination modes for carboxylate headgroups with lanthanide cations (La3+ and Y3+) have been identified at the air/water interface. Furthermore, the relative ratio between the different complexes was found to depend on the salt concentration in the subphase. Thus, the intricacies of the lanthanide complexation behavior impose structural complications on the self-assembly processes in the fatty acid monolayers. The underlying mechanisms regulating these processes are the main focus of our studies.
@@ 2.2 Brewster angle microscopy
We performed BAM measurements of the arachidic acid monolayer formed on Ce3+ solution to visualize the changes in morphological properties at different monolayer phases. The representative BAM images are shown in Fig. 2. As can be seen, the BAM data obtained for the gas and solid phases indicate the formation of a uniform sheet-like layer. This kind of BAM images are usually recorded for fatty acid formed on a subphase with metal ions [17].
Effects in the phase behavior of the arachidic acid monolayer induced by Ce were observed in BAM measurements in the collapse region (see Fig. 2). We performed BAM studies of the arachidic acid monolayer at two different temperatures, C and C. BAM images taken at C showed the formation of large 3D crystalline-like aggregates. Such disordered structures are often observed in BAM studies of the collapse behavior of fatty acid monolayers. In contrast, the post-collapse BAM data obtained for the low temperature monolayer (C) revealed the mosaic-like texture of the monolayer. As shown in Figs. 2(e), 2(f), the monolayer surface is divided into many two-dimensional patches, which can be identified as anisotropic domains that appear homogeneously across the monolayer upon compression. It should be emphasized that no growth of these domains into crystalline-like aggregates was observed upon further compression of the monolayer. In addition, BAM images taken at C near the collapse pressure revealed another interesting feature of the deformation processes in the collapse. As can be seen in Fig. 2(e), the surface of the monolayer is weakly grooved in the direction parallel to the barrier, which is a clear indication of the preferred deformation direction at the beginning of the collapse.
Taken together, the presented BAM studies of arachidic acid monolayers formed on a subphase containing Ce3+, revealed differences in the morphological properties of two observed collapsed modes. Thus, in our further discussion, the collapse with disordered morphology will be referred to as collapse mode 1; whereas the collapse with specific mosaic-like morphology will be referred to as collapse mode 2.
Note that similar results are reported in [19], where two qualitatively different collapse morphologies are also observed in fatty acid monolayers. They used phase contrast microscopy to study the collapse processes in the stearic acid monolayer as a function of the pH of the water subphase. According to the experimental data [19], the occurrence of collapse structures seen as bright random cracks was detected at , while a surface roughening collapse pattern was observed at .
@@ 2.3 Grazing incidence diffraction
Five series of GID measurements were performed in our studies. Each series was designed to study the structure of the arachidic acid monolayer at a different combination of surface pressure and temperature. There is one control series in which the arachidic acid monolayers were deposited on pure water subphase and four series in which arachidic acid monolayers were formed on a 20 mM Ce(NO3)6H2O salt solution. Other series were designed to study the solid state and the post-collapse state of the monolayer in the presence of Ce3+ ions in the liquid subphase. Within each series, except the control, multiple samples of the monolayer were prepared and measured. This was done to ensure reproducibility. The monolayers were compressed to a specified surface pressure at a specified temperature chosen for each series. The results are shown in Fig. 3. The measurements from the most representative samples are selected to be shown. From these selected samples, all collected data were integrated into a single GID pattern to represent each series. Some measurements were taken with different resolution or different scan area. These measurements are superimposed on the GID pattern, which is indicated by vertical dashed lines. It is important to note that for some measurements the GID pattern has evolved over time. Since the data is integrated over all measurements, different stages of structural change are captured as "motion blur" in Fig. 3. For such cases, the time evolution is shown in Figs. 4, 5 and 6.
@@
An arachidic acid monolayer was formed on the surface of pure water and compressed to a surface pressure mN/m at temperature . In this series of measurements [see Fig. 3(a)] a typical GID pattern for a Langmuir monolayer was observed: a triplet of peaks , , and . The brighter peak at the horizon of the monolayer is the peak, the peak above the horizon is a degenerate peak, in which two peaks and are superimposed. The arrangement of the peaks on top of each other indicates a hexagonal lattice: , . The tails of the fatty molecules are inclined with respect to the surface and are tilted in the direction of the nearest neighbor which corresponds to the NN tilt phase mode. We infer this from the arrangement of peaks on the horizon and directly above, which is characteristic of the NN tilt [5].
@@
In addition to a conventional GID pattern for a fatty acid monolayers in range C, we observed additional peaks in range A in this series. In range C we again see a peak at the horizon and a degenerated peak , above the horizon. However, the GID pattern observed in the control series, the peak is shifted horizontally to higher frequencies. This implies a lattice with tighter packing, compared to that of the arachidic monolayer on pure water: the surface area occupied by a molecule is reduced; 20 Å2 for series 1 versus 20.4 Å2 for control. The lattice is not hexagonal (), however the lattice angle is the same: . The structure has NN tilt.
Of particular interest are the bright peaks in the range A. We attribute these peaks to diffraction at a layer formed by Ce3+ ions adsorbed at the bottom of the monolayer on the surface of the subphase. This assumption is confirmed by two observations. First, the low spatial frequencies of the peaks imply a larger spatial period compared to the organic lattice. The coordinates of these peaks correspond to the fractional indices and of the reciprocal unit cell of the arachidic acid monolayer. This implies that Ce3+ ions form a superstructure underneath the arachidic acid monolayer. The unit cell of such a superstructure has a size of unit cells of the arachidic acid lattice. Second, the vertical size of the inorganic lattice peaks at ( Å-1) is significantly larger than that of the organic lattice peaks ( Å-1), indicating a greater vertical spatial localization of Ce3+ ions compared to that of the arachidic acid molecules. The formation of the Ce3+ layer at the subphase interface below the monolayer is also supported by the XSW measurements, as will be discussed in the next section. We will refer to the superstructure of Ce3+ ions as the inorganic superstructure.
@@
The GID pattern changed during the measurements, its evolution is shown in Fig. 4. Similar to series 1, it exhibits a typical GID pattern, characteristic of an NN tilt, in the range C [see Fig. 4(a)]. This pattern was first observed 3 hours after compression. The full range measurement shown in Fig. 3(c) was performed over the next 4 hours. A part of this measurement is also shown in Fig. 4(b) to emphasize that no peaks were detected in the B region during this period. The GID pattern of the inorganic superstructure was also observed [see peaks in range A in Fig. 3(c)]. The change in GID pattern was first observed 11 hours after compression [see Fig. 4(c)]. It evolved to a final GID pattern [see Fig. 4(d)] during another 7 hours of measurements. In this pattern we observed five peaks in the range C. Three more peaks appear in addition to the NN pattern. We interpret the pattern in Fig. 4(d) as diffraction from two coexisting packing modes. The additional peaks are an another triplet of , and , corresponding to a new packing mode. Since we can distinguish all three peaks at different , the lattice in this mode should have an oblique unit cell. We can also conclude that there is no tilt of the arachidic acid molecules, since all three peaks are on the horizon of the sample.
@@
Finally, 22 hours after the compression, we detected another peak at Å-1 in range B. We ascribe this peak with diffraction on a superstructure corresponding to the lattice of a new packing mode. The peak is identified by the fractional indices . In contrast to the diffraction on the inorganic superstructure, this peak is considerably smaller in the direction with Å-1. On the other hand, it is comparable in to the peaks of the organic structure in range C. Therefore, this is not a diffraction at the Ce3+ lattice, but diffraction on a superstructure composed of the arachidic acid molecules – organic superstructure. It can appear in the Langmuir monolayers due to the azimuthal ordering of the hydrocarbon chains of the molecules [21, 22]. Molecules of the monolayer can assume a denser packing due to a specific mutual orientation of the hydrocarbon chains of neighboring molecules, that is called herringbone (HB) an ordering when the azimuth angle between the hydrocarbon chains is 90∘, and pseudoherringbone (PHB) ordering when this angle is about 40∘ [23]. Similar organic superstructure peaks in the GID have also been reported in arachidic acid films in [24], but there the Langmuir-Blodgett multilayers on solid substrate were studied, not the monolayer on liquid. They report a superstructure with cell and an HB arrangement. The peak in our study can be ascribed to a superstructure and a PHB arrangement.
@@
To summarize the series 2, the monolayer was initially observed in a solid state phase with rectangular lattice, NN tilt and inorganic superstructure, identical to series 1. However, in contrast to series 1, a new packing mode was observed 11 hours after compression. We attribute this to the PHB ordering. This packing mode coexists with an initial phase.
@@
The monolayer samples in this series have been compressed beyond the collapse point. An integrated GID pattern is shown in Fig. 3(d). GID measurements for this sample were taken over a 41-hour period. The working temperature is C. The surface pressure of 52 mN/m was reached during compression, after which the surface area was kept constant throughout the measurements.
In the first GID pattern measured after compression, we again see a GID pattern characteristic of the NN tilt mode [see Fig. 5(a)]. The GID pattern of the inorganic superstructure was also detected after compression, as seen in the integrated pattern in Fig. 3(d): note the peak in the range A. However, the GID pattern has started to degrade, as shown in Fig. 5(b) measured 19 hours after compression. In addition, the peak from the inorganic superstructure is still visible 24 hours after compression. In the measurement 5 hours later, the peak from the inorganic superstructure has disappeared. Subsequently, the GID pattern of the monolayer in the range C also degraded, as shown in Fig. 5(e).
For the inorganic superstructure, only the peak was followed to save measurement time for resolving the structural change of the monolayer. However, it can be argued that should exist. Indeed, in this series the patterns characteristic of the NN tilt mode and the inorganic structure are detected, as in series 1 and at the beginning of series 2. Since it is attributed to an inorganic superstructure, the is also very likely to exist.
Thus, in series 3, the monolayer was compressed past the collapse point. Initially, the structure of the monolayer was similar to that of a solid phase monolayer as in series 1 and 2, but the structure began to degrade, and 41 hours after compression, only the peak corresponding to the short-range order was present, indicating the lateral disarrangement of the monolayer. The inorganic superstructure has also degraded as a result of the collapse.
@@
In this series, a new collapse mode of the arachidic acid monolayer was studied. The integrated GID pattern is shown in Fig. 3(e). The monolayer has been compressed beyond the collapse point at mN/m, after which the compression was stopped. The monolayer is studied at temperature C.
@@
The GID pattern in Fig. 3(e) is unusual for diffraction on a 2D polycrystalline system: the Bragg rods are rounded. The rounding of the peaks is exactly on the circle in reciprocal space and around its origin. In this sense, these peaks are similar to the Scherrer rings characteristic of diffraction on a 3D polycrystalline material. In our case, these diffraction rings are not closed. One can see the rounded rods instead. This suggests that collapse mode 2 is an intermediate state of the monolayer between a 3D and a 2D powder. In fact, the Langmuir film is still in a monolayer state because we still observe the GID pattern characteristic of a monolayer, namely a monolayer with PHB hydrocarbon chain ordering. Moreover, the positions of peaks , and are the same as in series 2 [see Figs. 3(c), 3(e) and Table 2.3]. Their coordinates are within 1% deviation between series 2 and 4. There is also a peak in range B [see Fig. 3(e)], again suggesting the PHB phase mode. However, these 2D domains are oriented in 3D with their normal deviating from the normal of the subphase. Presumably, in the collapse mode 2, the structure can be ascribed to a corrugated monolayer, as the domains are randomly oriented with a preferential orientation in the subphase plane.

### v2

@@ Abstract
Ce-induced effects on the self-assembly of arachidic acid monolayer was studied. The monolayers were formed on the cerium nitrate solution and studied in condensed and collapsed phases at two different temperatures and . Grazing incidence X-ray diffraction and X-ray standing waves were applied to monitor in real time the changes in molecular packing. A new type of molecular organization was found, in which the monolayer maintains its lattice structure despite being overcompressed. Instead of forming crystalline aggregates or anisotropic cracking, the collapsed monolayer at appeared to be corrugated. The diffraction pattern for the monolayer in a new collapse mode is represented by the unclosed diffraction rings with maxima near the sample horizon. Theoretical formalism based on the distorted wave Born approximation was developed for quantitative analysis of the full 2D maps of diffraction scattering. The collapsed monolayer at was shown to consist of crystalline flat domains, which are inclined from the horizontal position. The crystalline structure of a domain was identified as the pseudoherringbone packing mode.
@@ 1 Introduction
Scientific interest in the self-assembly of amphiphilic trivalent lanthanide complexes is motivated by their photophysical properties, which can be exploited for molecular sensing [1]. The practical realization of such systems requires the assembly of the lanthanide complexes into nanostructures [2]. One of the possible ways to create functional nanomaterials with predefined molecular architecture is the Langmuir-Blodgett technique [3], which involves the self-assembly of amphiphiles at the air-liquid interface into a ordered monolayer and subsequent transfer to a solid substrate. Charged headgroups of amphiphilic molecules are known to interact with metal ions when present in the liquid subphase [4]. In this way, structural organization of the monolayer defines the performance of a nanodevice. Self-assembly is driven not only by the sort of amphiphiles and the thermodynamic conditions under which the monolayer is formed [5], but can also be directed by interactions with ions in the subphase [6]. Therefore, further development of functional devices requires extensive and detailed studies of the structural organization of amphiphilic trivalent lanthanide complexes at liquid interfaces.
@@
There are two known scenarios of how the interaction with metal ions affects the structure of the Langmuir monolayer, as reported in [7] using divalent metal ions as an example. The first, more common scenario is that the interaction between headgroups and metal ions leads to the formation of a disordered thin monolayer of metal ions underneath the organic monolayer, which affects the compressibility of the monolayer, resulting in structures similar to high-pressure solid-phase monolayers [7]. The second scenario is the formation of the ionic monolayer ordered into a superstructure commensurate with the organic monolayer, which was first reported in [8]. The study reports the formation of a Cd2+ superstructure underneath an arachidic acid monolayer. The structure has a cell commensurate with the organic monolayer structure. The transition between these two scenarios of interaction of the organic monolayer with metal ions is known to be dependent on the pH of the subphase and the ion concentration. In [9], the interaction of behenic acid with a variety of divalent ions was studied, and an ion concentration threshold was observed. Above this threshold, the ions formed a superlattice, and below it the ion layer was disordered. Later studies [10, 11] report on forming ordered monolayers of lanthanide ions. They tested a wide range of trivalent lanthanides. They found that the ordering depends not only on the molecules forming the Langmuir monolayer, but also on the atomic number of a lanthanide. In addition, the crystalline structure of the ionic monolayer can be commensurate or incommensurate with the lattice of the organic molecules.
@@
Another interesting aspect is the collapse of Langmuir monolayers in the presence of metal ions. The study in [12] reports the spontaneous collapse of fatty acid monolayers in the presence of Ca2+ ions. It has been shown both experimentally [13] and theoretically [14] that the presence of Ca2+ ions changes the structure of the collapsed fatty acid monolayer. Even more interesting is that the collapsed structures show a high degree of order, namely the formation of inverted bi-layers. The formation of such inverted bi-layers is also reported for lanthanides [15]. Yet again, the structure of the collapsed layers is ion-specific.
@@
Thus, the interaction of fatty acids with lanthanide ions has shown complex and widely varying results. However, understanding the mechanisms of ion monolayer formation and its effect on collapse is relevant to the problems of functionalizing lanthanides into self-assembled nanostructures. This article focuses on the self-assembly of arachidic acid (AA) monolayers in the presence of Ce3+ ions and their structural reorganization while compressed beyond the collapse point. We have studied the structure of fatty acid monolayers interacting with metal ions at the air-liquid interface at different surface pressures and temperatures. As the primary tool for monitoring the monolayer structure, we used grazing incidence X-ray diffraction (GID) at the bending magnet synchrotron radiation source. This choice is due to the sensitivity of GID to both translational and orientational lateral ordering of the monolayers. In addition, simultaneously with GID experiments, we performed the X-ray standing wave (XSW) measurements to monitor the vertical localization of Ce3+ ions during the structural reorganization at the collapse.
@@ 2.1 Surface pressure-area isotherms
The effect of Ce3+ ions on the thermodynamic properties of the AA monolayer was investigated in our additional series of experiments using compression isotherm measurements. We have studied the behavior of the AA monolayer, formed on aqeous solution of Ce(NO3)6H2O salt and on pure water as a control. All measurements were performed at a temperature of C. The typical - isotherms are shown in Fig. 1. Based on these experimental results, the key feature of the AA monolayer formed in the presence of Ce3+ is the absence of a liquid phase, as evidenced by a high monolayer compressibility – 2.2 m/mN.
Our observations are consistent with the results reported in [16], where detailed studies of the binding interaction of lanthanide ions (La3+ and Y3+) with AA monolayers at submicromolar salt concentrations were reported. Note that similar changes in the isotherms were observed in [17], where the effect of Cu2+ on the organization and morphology of the AA monolayer was examined.
Generally, metal binding with carboxylate acid groups is recognized to promote the formation of a rigid monolayer with higher crystallinity and reduced molecular area [18]. The interaction of divalent and trivalent metal ions with carboxylic groups of fatty acids is strongly dependent on such parameters as solution pH, temperature, salt concentration. The type of metal-carboxylate complex (ionic, unidentate, chelating bridging bidentate, polymeric tridentate) is determined to a large extent by metal ion specific nature, i.e. ion size, coordination structure, oxidation state in solution, etc. Large Ln3+ ions with high coordination numbers exhibit rather intricate binding behavior [16], resulting in a wide variety of the lanthanide-carboxylate binding interactions. According to the data reported in [16], at least three different coordination modes for carboxylate headgroups with lanthanide cations (La3+ and Y3+) have been identified at the air/water interface. Furthermore, the relative ratio between the different complexes was found to depend on the salt concentration in the subphase. Thus, the intricacies of the lanthanide complexation properties impose structural complications of the self-assembly processes in the fatty acid monolayers. The underlying mechanisms regulating these processes are the main focus of our studies.
@@ 2.2 Brewster angle microscopy
We performed BAM measurements of the AA monolayer formed on Ce3+ solution to visualize the changes in morphological properties at different monolayer phases. The representative BAM images are shown in Fig. 2. As can be seen, the BAM data obtained for the gas and solid phases indicate the formation of a uniform sheet-like layer. This kind of BAM images are usually recorded for fatty acid formed on a subphase with metal ions [17].
Ce-induced effects in the phase behavior of the AA monolayer were observed in BAM measurements in the collapse region (see Fig. 2). We performed BAM studies of the AA monolayer at two different temperatures, C and C. BAM images taken at C showed the formation of large 3D crystalline-like aggregates. Such disordered structures are often observed in BAM studies of the collapse behavior of fatty acid monolayers. In contrast, the collapse BAM data obtained for the low temperature monolayer (C) revealed the mosaic-like texture of the monolayer. As shown in Figs. 2(e), 2(f), the monolayer surface is divided into many patches, which can be identified as anisotropic domains that appear homogeneously across the monolayer upon compression. It should be emphasized that no growth of these domains into crystalline-like aggregates was observed upon further compression of the monolayer.
Taken together, the presented BAM studies of AA monolayers formed on a subphase containing Ce3+, revealed differences in the morphological properties of two observed collapsed modes. Thus, in our further discussion, the collapse with disordered morphology will be referred to as collapse mode 1; whereas the collapse with specific mosaic-like morphology will be referred to as collapse mode 2.
Note that similar results are reported in [19], where two qualitatively different collapse morphologies are also observed in fatty acid monolayers. They used phase contrast microscopy to study the collapse processes in the stearic acid monolayer as a function of the pH of the water subphase. According to the experimental data obtained in [19], the occurrence of collapse structures seen as bright random cracks was detected at , while at a surface roughening collapse pattern was observed.
@@ 2.3 Grazing incidence diffraction
Five series of GID measurements were performed in our studies. Each series was designed to study the structure of the AA monolayer at different surface pressure and temperature. There is one control series in which the AA monolayers were deposited on pure water subphase and four series in which AA monolayers were formed on a 20 mM Ce(NO3)6H2O salt solution. Within each series, except the control, multiple samples of the monolayer were prepared and measured. This was done to ensure reproducibility. The results are shown in Fig. 3. The measurements from the most representative samples are selected to be shown. From these selected samples, all collected data were integrated into a single GID pattern to represent each series. Some measurements were taken with different resolution or different scan area. These measurements are superimposed on the GID pattern, which is indicated by vertical dashed lines. It is important to note that for some measurements the GID pattern has evolved over time. Since the data is integrated over all measurements, different stages of structural change are captured as "motion blur" in Fig. 3. For such cases, the time evolution is shown in Figs. 4, 5 and 6.
@@
An AA monolayer was formed on the surface of pure water and compressed to a surface pressure mN/m at temperature . In this series of measurements [see Fig. 3(a)] a typical GID pattern for a Langmuir monolayer was observed: a triplet of peaks , , and . The brighter peak at the horizon of the monolayer is the peak, the peak above the horizon is a degenerate peak, in which two peaks and are superimposed. The arrangement of the peaks on top of each other indicates a hexagonal lattice: , . The tails of the fatty molecules are inclined with respect to the surface and are tilted in the direction of the nearest neighbor which corresponds to the NN tilt phase mode. We infer this from the arrangement of peaks on the horizon and directly above, which is characteristic of the NN tilt [5].
@@
In range C we again see a peak at the horizon and a degenerated peak , above the horizon. However, the GID pattern observed in the control series, the peak is shifted horizontally to higher frequencies. This implies a lattice with tighter packing, compared to that of the AA monolayer on pure water: the surface area occupied by a molecule is reduced; 20 Å2 for series 1 versus 20.4 Å2 for control. The lattice is not hexagonal (), however the lattice angle is the same: . The structure has NN tilt.
Of particular interest are the bright peaks in the range A. We attribute these peaks to diffraction at a layer formed by Ce3+ ions adsorbed at the bottom of the monolayer on the surface of the subphase. This assumption is confirmed by two observations. First, the low spatial frequencies of the peaks imply a larger spatial period compared to the organic lattice. The coordinates of these peaks correspond to the fractional indices and of the reciprocal unit cell of the AA monolayer. This implies that Ce3+ ions form a superstructure underneath the AA monolayer. The unit cell of such a superstructure has a size of unit cells of the AA molecular lattice. Second, the vertical size of the inorganic lattice peaks at ( Å-1) is significantly larger than that of the organic lattice peaks ( Å-1), indicating a greater vertical spatial localization of Ce3+ ions compared to that of the AA molecules. The formation of the Ce3+ layer at the subphase interface below the monolayer is also supported by the XSW measurements, as will be discussed in the next section. We will refer to the superstructure of Ce3+ ions as the inorganic superstructure.
@@
The GID pattern changed during the measurements, its evolution is shown in Fig. 4. Similar to series 1, it exhibits a typical GID pattern, characteristic of an NN tilt, in the range C [see Fig. 4(a)]. This pattern was first observed 3 hours after compression. The full range measurement shown in Fig. 3(c) was performed over the next 4 hours. A part of this measurement is also shown in Fig. 4(b) to emphasize that no peaks were detected in the B region during this period. The GID pattern of the inorganic superstructure was also observed [see peaks in range A in Fig. 3(c)]. The change in GID pattern was first observed 11 hours after compression [see Fig. 4(c)]. It evolved to a final GID pattern [see Fig. 4(d)] during another 7 hours of measurements. In this pattern we observed five peaks in the range C. Three more peaks appear in addition to the NN pattern. We interpret the pattern in Fig. 4(d) as diffraction from two coexisting packing modes. The additional peaks are another triplet of , and , corresponding to a new packing mode. Since we can distinguish all three peaks at different , the lattice in this mode should have an oblique unit cell. We can also conclude that there is no tilt of the AA molecules, since all three peaks are on the horizon of the sample.
@@
Finally, 22 hours after the compression, we detected another peak at Å-1 in range B. We ascribe this peak with diffraction on a superstructure corresponding to the lattice of a new packing mode. The peak is identified by the fractional indices . In contrast to the diffraction on the inorganic superstructure, this peak is considerably smaller in the direction with Å-1. On the other hand, it is comparable in to the peaks of the organic structure in range C. Therefore, this is not a diffraction at the Ce3+ lattice, but diffraction on a superstructure composed of the AA molecules – organic superstructure. It can appear in the Langmuir monolayers due to the azimuthal ordering of the hydrocarbon chains of the molecules [21, 22]. Molecules of the monolayer can assume a denser packing due to a specific mutual orientation of the hydrocarbon chains of neighboring molecules, that is called herringbone (HB) packing – an ordering when the azimuth angle between the hydrocarbon chains is 90∘, and pseudoherringbone (PHB) packing when this angle is about 40∘ [23]. Similar organic superstructure peaks in the GID have also been reported for AA films in [24], but there the Langmuir-Blodgett multilayers on solid substrate were studied. They report a superstructure with cell and an HB arrangement. The peak in our study can be ascribed to a superstructure and a PHB arrangement. It should be emphasized that in our experiments we were able to monitor in real time the gradual transition from the NN tilt mode to PHB packing.
@@
To summarize the series 2, the monolayer was initially observed in a solid phase with rectangular lattice, NN tilt and inorganic superstructure, identical to series 1. However, in contrast to series 1, a new packing mode was observed 11 hours after compression. We attribute this to the PHB ordering. This packing mode coexists with an initial phase.
@@
The monolayer samples in this series have been compressed beyond the collapse point. An integrated GID pattern is shown in Fig. 3(d). GID measurements for this sample were taken over a 41-hour period. The temperature is C. The surface pressure mN/m was reached during compression, after which the surface area was kept constant throughout the measurements.
In the first GID pattern measured just after compression, we again see a GID pattern characteristic of the NN tilt mode [see Fig. 5(a)]. The GID pattern of the inorganic superstructure was also detected after compression, as seen in the integrated pattern in Fig. 3(d): note the peak in the range A. However, the GID pattern has started to degrade, [Fig. 5(b)] measured 19 hours after compression. In addition, the peak from the inorganic superstructure is still visible 24 hours after compression. In the measurement 5 hours later, the peak from the inorganic superstructure has disappeared. Subsequently, the GID pattern of the monolayer in the range C also degraded, as seen in Fig. 5(e).
For the inorganic superstructure, only the peak was monitored to save measurement time for resolving the structural change of the monolayer. Since it is attributed to an inorganic superstructure, the is also very likely to exist.
Thus, in series 3, the monolayer was compressed beyond the collapse point. Initially, the structure of the monolayer was similar to that of a solid phase monolayer as in series 1 and 2, but the structure began to degrade, and 41 hours after compression, only the peak corresponding to the short-range order was present, indicating the lateral disordering of the monolayer. The inorganic superstructure has also degraded as a result of the collapse.
@@
In this series, a new collapse mode of the AA monolayer was studied. The integrated GID pattern is shown in Fig. 3(e). The monolayer has been compressed beyond the collapse point at mN/m, after which the compression was stopped. The monolayer was studied at temperature C.
@@
The GID pattern in Fig. 3(e) is unusual for diffraction on a 2D polycrystalline system: the Bragg rods are rounded. The rounding of the peaks is exactly on the circle in reciprocal space and around its origin. In this sense, these peaks are similar to the Scherrer rings characteristic of diffraction on a 3D polycrystalline material. In our case, these diffraction rings are not closed. One can see the rounded rods instead. This suggests that collapse mode 2 is an intermediate state of the monolayer between a 3D and a 2D powder. In fact, the Langmuir film retains high degree of structural organization because we still observe the GID pattern characteristic of an ordered monolayer, namely a monolayer with PHB hydrocarbon chain ordering. Moreover, the positions of peaks , and are the same as in series 2 [see Figs. 3(c), 3(e) and Table 2.3]. Their coordinates are within 1% deviation between series 2 and 4. There is also a peak in range B [see Fig. 3(e)], again suggesting the PHB packing mode. However, these 2D domains are oriented in 3D with their normal deviating from the normal of the subphase. Presumably, in the collapse mode 2, the structure can be ascribed to a corrugated monolayer, as the domains are randomly oriented with a preferential orientation in the subphase plane.